Jill pushes a lawn mower at an angle of 45o measured from the horizontal. If she applies an average force of 98.5 N for a distance of 2.50 meters, how much work did she do on the lawn mower?
1
Expert's answer
2020-11-06T10:14:34-0500

By the definition of the work done, we have:


"W=Fscos\theta=98.5\ N\cdot 2.5\ m\cdot cos45^{\circ}=174.12\ J."

Answer:

"W=174.12\ J."